# Brain Tumor Classification on MRI in Light of Molecular Markers

### v3

@@ 4 Discussion
We present a viable CNN-based technique for predicting deletions of the 1p/19q chromosomal arm. A significant challenge in applying deep learning algorithms to medical imaging is the lack of sufficient datasets. For medical diagnosis, current transfer learning methods rely on various publicly available models trained on large ImageNet datasets. However, these models often generate a substantial number of medical images, which can compromise the accuracy of clinical diagnosis [16].
@@
A small dataset [18] can result in substantial training errors. In order to solve this problem, Gaussian noise is added during the training phase to improve the model’s normalization ability and fault tolerance. Adding noise prevents the network from memorizing the training samples, as they are constantly changing. As a result, the network weights are reduced, making the network more robust and leading to a lower generalization error. Since fresh samples are drawn close to existing samples in the input space, the shape of the input space is smoothed out. This makes it easier for the network to learn the mapping function, resulting in faster and more effective learning. In the next step, we will explore using diffusion [20] models to generate small samples, combine them with LLM for 3D reconstruction [31, 32, 33], and apply compression [21, 22, 23] methods for deployment on embedded devices [24, 25, 26, 29, 30].

### v4

@@ 4 Discussion
We present a viable CNN-based technique for predicting deletions of the 1p/19q chromosomal arm. A significant challenge in applying deep learning algorithms to medical imaging is the lack of sufficient datasets. For medical diagnosis, current transfer learning methods rely on various publicly available models trained on large ImageNet datasets. However, these models often generate a substantial number of medical images, which can compromise the accuracy of clinical diagnosis [10, 25].
@@
A small dataset [27] can result in substantial training errors. In order to solve this problem, Gaussian noise is added during the training phase to improve the model’s normalization ability and fault tolerance. Adding noise prevents the network from memorizing the training samples, as they are constantly changing. As a result, the network weights are reduced, making the network more robust and leading to a lower generalization error. Since fresh samples are drawn close to existing samples in the input space, the shape of the input space is smoothed out. This makes it easier for the network to learn the mapping function, resulting in faster and more effective learning. In the next step, we will explore using diffusion [29] models to generate small samples, combine them with LLM for 3D reconstruction [36, 37, 38], and apply compression [2, 7, 30, 31] methods for deployment on embedded devices [9, 32, 33, 3, 34, 35].